11. A test charge produced an electric field, E, at point 3 m away from the charge. The point where the field is half its original values is located at A. 1.5 m from the charge C. 4.5 m from the charge B. 6 m from the charge D. at the point of charge 12. The electric field of a body is directly related to its ________. A. momentum B. kinetic energy C. potential energy D. charge it carries 13. Charge q1 (26.0 µC) and q2 (-47.0 µC) experienced an electrostatic force of 5.70 N between them at a distance of ____ A.1.39 m B.1.93 m C.2.14 x 10-19 D.1.464 x 10-8


1
Expert's answer
2022-02-02T12:48:31-0500

11.

"E=\dfrac{kq}{3^2}→r=\sqrt\dfrac{2kq}{E}=3\sqrt\dfrac{2kq}{kq}=4.2\text{ m}."


12. D. 

13.

"r=\sqrt{kq_1q_2/F}=1.39\text{ m}."